40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 25 03:14:46 2014, ericq, Update, LSC, locking activity 
    Reply  Fri Apr 25 13:18:08 2014, Dark Jamie, Update, LSC, locking activity m10008_f1_bg.jpg
Message ID: 9853     Entry time: Fri Apr 25 03:14:46 2014     Reply to this: 9855
Author: ericq 
Type: Update 
Category: LSC 
Subject: locking activity 

[ericq, Jenne, Zach]

We spent some time tonight trying to push our CARM locking further, to little avail. DARM/CARM loop oscillations kept sneaking up on us. We measured some MC2 motion -> REFL11 Transfer Functions to see if we could see CARM plant features; plots will come in the near future...

ELOG V3.1.3-